OpenStudy (anonymous):

A yardstick measures 1/4 by 3 by 36 inches. How many yard sticks will fit in a box 3 inches wide and 36 inches high, if the girth of the box is 30 inches?

OpenStudy (radar):

Girth = 30 inches Girth =3 + 3 + x + x 30 = 6 +2x 2x=30-6=24 x=12 yardsticks are 1/4 in thick 12/.25=48 48 yardsticks could be packed in that box.

OpenStudy (anonymous):

thanx

OpenStudy (radar):

you're welcome